Convert the given lengths from the derived units to meters. 0.425 mm m 859 Mm m I 33.5 km
Answer #1

A. 1 mm = 10-3 m

So 0.425 mm = 4.25 × 10-4 m

B. 1 Mm = 106m

So 859 Mm= 859 × 106 m = 8.59 × 108 m

C. 1 km = 103 m

So 33.5 km = 3.35 × 104 m

Just multiply them with their conversion factor and then present them into scientific notation